OpenStudy (anonymous):

Simplify. Express the product as a rational exponent (x^3/4)^3/7

OpenStudy (anonymous):

when you have a power to a power, you multiply the powers example (x^2)^3=x^2*x^2*x^2=x^6 2*3=6 so, start by multiplying 3/4 and 3/7

OpenStudy (anonymous):

top & bottom?

OpenStudy (anonymous):

3 x 3 = 9

OpenStudy (anonymous):

i got my answer!!!!! :))))))))

OpenStudy (anonymous):

3x3=9, 4x7=28 the answer would be 28th root of x^9 or x^(9/28) depending on the form

OpenStudy (anonymous):

I got it on my own!!!!!!!!!!!!!!! :))))))))))Thank you!!!!!!!!!!!!!!!!!!!!1